OpenStudy (sf_r):

Factor completely with showing all work x^2-16x+64

OpenStudy (sf_r):

-8 and -8

OpenStudy (xapproachesinfinity):

that's a perfect square

OpenStudy (sf_r):

64 and x^2?

OpenStudy (xapproachesinfinity):

(a-b)^2=a^2-2ab+b^2

OpenStudy (xapproachesinfinity):

like this find a and b in your case

OpenStudy (sf_r):

a is x^2 and b is -16x

OpenStudy (xapproachesinfinity):

b=8 not -16x

OpenStudy (xapproachesinfinity):

64=8^2

OpenStudy (sf_r):

I am confused!

OpenStudy (xapproachesinfinity):

if a=x and b=8 then you have x^2-2(8)x+8^2=(x-8)^2

OpenStudy (sf_r):

I mostly do the Asterisk method probably thats why

OpenStudy (sf_r):

use the*

OpenStudy (xapproachesinfinity):

no need for this one

OpenStudy (xapproachesinfinity):

this one is the easiest of those factoring

OpenStudy (sf_r):

@satellite73

OpenStudy (sf_r):

@ILovePuppiesLol

OpenStudy (sf_r):

@Kainui